Here is one that I am struggling with the directions.
Find the polynomial for the perimeter and the area. Do not factor.
there is a rectangle with the length labeled z+5 and the width is labeled z.
I think the perimeter would be 2z+2(z+5) and the area may be z(z+5). Thanks

Answer by stanbon(75887)

Find the polynomial for the perimeter and the area. Do not factor.
there is a rectangle with the length labeled z+5 and the width is labeled z.
--------------------------
Perimeter = 2(length + width)
P = 2(z+5 + z) = 2(2z + 5) = 4z + 10
-----------------------------------------
Area = width * length
A = z(z+5) = z^2 + 5z
=============================
Cheers,
Stan H.